OpenStudy (anonymous):

Find the surface area of an orange with a 3'' radius.

OpenStudy (anonymous):

24sq. in. 36sq. in. 144sq. in.

OpenStudy (anonymous):

\[4 \pi r ^{2}\] So if r=3: \[4 \pi (3)^{2}= 36 \pi\] is your answer, i didn't mention the unit because its not given.

OpenStudy (anonymous):

36 sq.inches.

OpenStudy (anonymous):

thank you